# Large Language Models Powered Context-aware Motion Prediction

### v1

@@ III-C LLM Augmented Motion Prediction
Performance comparison. We compared our LLM augmented motion prediction model with MTR on the validation set of WOMD. We trained both models on 20% WOMD dataset. Our main results are in Table III. V, P, C are abbreviations of vehicle, pedestrian, cyclist, respectively. The result shows that our LLM augmented motion prediction model (with generated similar context information) outperformed MTR among all three agents (i.e. vehicle, pedestrian, and cyclist), and the average mAP increased by 0.95% compared with MTR.
@@
Performance Comparison. With LLM-augmented traffic context information, the motion prediction accuracy of different traffic agents improved. The best scores are expressed in bold.
Ablation Study. This study evaluates the contribution of each component within the transportation context generated by the LLM toward enhancing motion prediction accuracy. An ablation experiment was conducted by randomly selecting 5% of scenarios (approximately 24,000 scenarios) from the WOMD training set. The evaluation involved comparing the baseline model against the LLM-augmented model on the minimal part of WOMD validation set with traffic context information generated by LLM. The results are presented in Table IV. I, A, and S are abbreviations of intention, affordance, and scenario, respectively. V-mAP, P-mAP, and C-mAP mean mAP scores for vehicle, pedestrian, and cyclist, respectively. The results highlight the incremental value added by incorporating various types of transportation context information. Specifically, augmenting the model with all three types of transportation context information yielded the highest mAP score. Compared to the baseline MTR model, the final LLM-augmented model demonstrated a significant improvement, with a 1.67% increase in the mAP score.
@@
Ablation Experiment. Results demonstrate that the model that uses three types of transportation context information gets the highest mAP score. The highest mAP score for each types of agent and average mAP score are expressed in bold.

### v3

@@ III-B LLM Augmented Motion Prediction
Performance comparison. We compared our LLM augmented motion prediction model with MTR on the validation set of WOMD. We trained both models on the 20% WOMD dataset. Our main results are in Table III. Our LLM-augmented model exhibited superior performance over the MTR model across all agent types (vehicles, pedestrians, cyclists) on the WOMD validation set. Notably, the average mAP saw an enhancement of 0.95%, underscoring the augmented model’s effectiveness.
@@
The best scores are expressed in bold.
Ablation Study. This study evaluates the contribution of each component within the transportation context generated by the LLM toward enhancing motion prediction accuracy. An ablation experiment was conducted by randomly selecting 5% of scenarios (approximately 24,000 scenarios) from the WOMD training set to train a model. The evaluation involved comparing the baseline model against the LLM-augmented model on the minimal part of the WOMD validation set with traffic context information generated by LLM. The results are presented in Table IV. I, A, and S are abbreviations of intention, affordance, and scenario, respectively. V-mAP, P-mAP, and C-mAP mean mAP scores for vehicle, pedestrian, and cyclist, respectively. The results highlight the incremental value added by incorporating various types of transportation context information. Incorporation of all context types resulted in the highest mAP scores, marking a 1.67% increase over the baseline MTR model, thereby validating the efficacy of LLM-augmented context in enhancing motion prediction.
@@
The highest mAP score for each type of agent and the average mAP score is expressed in bold.